OpenStudy (anonymous):

lisa bought 15 cds in december .of those 3 were jazz ,the rest were world music.How many world music discs did she buy ?

OpenStudy (anonymous):

15 -3 = 12

OpenStudy (anonymous):

wow a nice question! 15-3 =12 there are 15 cds out of which 3 were of jazz so 15-3 =12 rest of them were music cds so 12 is the answer!

OpenStudy (jebonna):